Find the vertex for the function f(x)=x^2+6x-4
a. 2,12
b. -3,-13
c. -1,-9
d. 0,-4

Answer by oscargut(686)

Find the vertex for the function f(x)=x^2+6x-4
In gral if f(x)=ax^2+bx+c then V=(-b/2a,f(-b/2a))
so in this case V=(-6/2,f(-6/2))=(-3,-13)
then option b)is the correct